CALCULUS POST-MORTEM: WHAT HAVE YOU LEARNED?

1. Calculations

Much of your homework consisted of doing countless calculations, mostly: • derivatives

• integrals (in 1 dimension)

• limits

• (some) differential equations You’ll get more of these in classes/applications; now you can do them, and often quickly.

2. Math expression

Other than doing calculations, a key goal has been for you to learn to express yourself mathematical, as you do with words: if you have an idea that can be formulated mathematical, can you express it, can you convey it to someone?

• writing math: proofs, rigorous argument

• reading math: seeing what’s happening in equations

• speaking math: being able to express an idea orally

3. Math understanding

You now have a better feel for functions and how they behave: rather than just shuffling symbols, you can look at some computations and equations and see if they’re correct.

You understand